# Global survey of star clusters in the Milky Way VIII. Cluster formation and evolution

### v1

@@ 5.2 Comparison of fit statistics
Secondly, we find that the parameters of violent relaxation are not significantly constrained beyond their priors. To quantify this, we consider the Kolmogorov-Smirnov statistic (KS-statistic) of the marginalised posteriors for the respective priors, that is, the maximum of the absolute difference between the cumulative distribution functions of the marginalised posterior and the prior. As we see in Table 6, the KS-statistic is significantly lower for $t_{\mathrm{V}}$ than for the parameters of the CIMF, independent of the fit statistic used. For $n_{\mathrm{b}}$, the KS-statistic is still below all but the lowest value found for any combination of fit statistic and CIMF parameter. The reason for this may be that while the entire cluster sample contains implicit information on the shape of the CIMF, only young clusters with $\tau<8$ contain information on violent relaxation, which is a smaller sample containing only 497 clusters. Moreover, this sample was already used to derive prior constraints, limiting the information that could be extracted additionally by the fitting procedure.
@@ 5.2.1 Construction of synthetic cluster samples
In order to better understand the differences between the fit statistics, we consider now synthetic cluster samples as realisations of a given model. We construct such a realisation by drawing a sufficient number $N$ of clusters from the model, applying random Gaussian errors according to the error model from Sect. 4.1, and then evenly distributing them across a heliocentric disc of radius $R=\sqrt{{N}/\left({\pi\Sigma_{\mathrm{tot}}}\right)}$ such that the cluster surface density corresponds to the observed value of $\Sigma_{\mathrm{tot}}=134.1\,\mathrm{kpc^{-2}}$. In analogy to the magnitude-dependent completeness limit, a completeness limit depending on the logarithm of its mass is drawn for each cluster. These limits are drawn from a Gaussian with mean and standard deviation of the form
@@
with values $w_{\mathrm{mean}}=0.60\,\mathrm{kpc},\,v_{\mathrm{mean}}=0.84\,\mathrm{kpc}$ for the mean and $w_{\mathrm{std}}=0.43\,\mathrm{kpc},\,v_{\mathrm{std}}=0.12\,\mathrm{kpc}$ for the standard deviation obtained by performing linear regression on the completeness limits of the observed cluster sample. A minimum completeness distance of 0.2 kpc was chosen to disallow divergent contributions to the synthetic CAMF. These completeness limits then also allow the definition of a sufficient number of clusters such that the radius of the resulting heliocentric disc is at least as large at the greatest completeness distance. In practice, this was done by choosing $N$ such that $R\geq 6.8\,\mathrm{kpc}$.
The final synthetic sample is then obtained by selecting only clusters within their completeness limits. While this procedure produces samples with similar properties as the observed one by coarsely approximating the measurement and selection process used to arrive at a completeness-corrected sample, it cannot directly fix global properties of the samples such as number of clusters and total cluster surface density, which as a result vary between realisations and models.
For each of the base, $\chi^{2}$-fit, $D_{\mathrm{KL}}$-fit, and $E[\ln P]$-fit models, 5000 realisations of synthetic cluster samples were drawn and the fit statistics with respect to the corresponding models computed.
While the total cluster surface density and cluster number of these realisations is not fixed, their construction causes them to reproduce the observed cluster surface density well, with a stochastic scatter of about 6 kpc-2. The corresponding number of clusters also varies, with systematic deviations from the cluster number of the observed sample being introduced by the differing structure of the model CAMFs. In particular, realisations of $\chi^{2}$-fit model and $E[\ln P]$-fit model tend to produce fewer clusters, while realisations of the $D_{\mathrm{KL}}$-fit model produce cluster numbers consistent with the observed sample.
@@ 5.2.2 Fit statistic distributions for model realisations
Further, the fit-statistic distributions allow us assign a $Q$-value to the observed value of the fit statistic. This represents the quantile at which the observation sits and gives the fraction of model realisations with a lower degree of agreement with the model as measured by the statistic, that is, a larger value for $\chi^{2}$ and KLD, and a smaller value for $E[\ln P]$. This $Q$-value gives an indication how typical the level of disagreement between model and data would be for a realisation of the model, both in terms of atypically high or low values of the statistic. A very low $Q$-value means that it is unlikely that a random realisation of the model will disagree with the model at least as much as the data. As such, it signals that the model does not fit the data well. Conversely, a very high $Q$-value means that it is unlikely for random realisations of the model to agree with the model better than the data. In this sense, it signals that the model may be ‘overfitted’. What we refer to here as overfitting is when a model is fitted to not just a data set’s general features, but also its specific noise. Such a model will produce spurious features, which is why this is an undesirable property. When overfitting occurs for a given combination of data and fit statistic, the data will match the model better than a typical realisation of the model itself, as a greater discrepancy is expected simply from the stochastic noise included in the forward modelling.
@@
In the $D_{\mathrm{KL}}$-fit model, the value of $\chi^{2}_{\mathrm{red}}$ falls within the central 68% (corresponding to a Gaussian $1\sigma$-interval), and the value of $D_{\mathrm{KL}}$ is still fairly typical with a $Q$-value at 14%. However, for $E[\ln P]$, we have a very low $Q$-value of 2.7%. As we will discuss in Sect. 5.2.3, this may indicate the model realisations being systematically more centrally concentrated than the observed sample. The $D_{\mathrm{KL}}$-fit model samples also reproduce the observed cluster number well.
@@ 5.2.3 General remarks and conclusions
From our analysis of the fit statistics of synthetic cluster sample, it is clear that one factor which limits parameter fitting is the large impact of sampling noise on the fit statistics, which limits their ability to discriminate between models based on the observed cluster sample. This is likely because our fit statistics, as we constructed them in Sect. 4.2, 4.3 and 4.4, do not take the Poisson noise of the cluster sample into account correctly. In our sample, faint clusters have large weight, which increases their impact on the noise.
Of our best-fitting models, the $D_{\mathrm{KL}}$-fit model performs the best across all fit statistics and reproduces the cluster number when sampled. It reproduces the observed sharp age limit while still allowing for high-age high-mass clusters with $\tau>8.5,\,\mu>3.5$ when accounting for uncertainties in mass determination. A caveat remains in that the synthetic sampling of the model shows that the value of $E[\ln P]$ for data and model is lower than for typical realisations. This may suggest that the model CAMF is more concentrated in the age-mass plane than the observed CAMF.
As such, we take the $D_{\mathrm{KL}}$-fit model to be the best model we found in our investigation of model parameter space.
@@ 5.3 Parameter correlations
Previously, we evaluated the uncertainties of the model parameters individually based on their marginalised one-dimensional posteriors. We now consider the full eleven-dimensional posterior to derive a better understanding of the uncertainties of the model parameters and their correlations.
As the high dimensionality of the parameter posterior distribution makes it challenging to analyse or visualise the full complexity of the posterior, we use principal component analysis in an attempt to make the posterior tractable without losing too much information.
This is done by first normalising the parameters to zero mean and unit standard deviation, and then performing an eigenvector decomposition on the covariance matrix of the posterior in normalised parameter space. The resulting eigenvectors $\mathsf{u}_{i}$ for indices $i=0\ldots 10$, ordered by the size of the corresponding eigenvalues, form an orthonormal basis of the normalised parameter space and are uncorrelated in the sense that $\mathrm{Cov}(\mathsf{u}_{i},\mathsf{u}_{j})=\lambda_{i}\delta_{ij}$ where $\lambda_{i}$ is the eigenvalue corresponding to $\mathsf{u}_{i}$. This means that the eigenvalues represent the variance of the posterior distribution along the direction of the corresponding eigenvectors. If we consider the distribution in the eigenvector basis, then the $i$-th parameter has zero mean, standard deviation $\sqrt{\lambda_{i}}$ and no linear correlation to another parameter. While higher-order correlations certainly exist, this at least allows us to consider independence of these principal components as a first-order approximation. In this picture, the posterior corresponds to an ellipsoidal distribution with principal axes given by the eigenvectors $\mathsf{u}_{i}$.
As such, if $\lambda_{i}$ is smaller, then the posterior is more constrained along the axis $\mathsf{u}_{i}$, while it is less constrained if $\lambda_{i}$ is larger, with the vectors with largest eigenvalues having the greatest contribution to the uncertainty of the parameters. Conversely, the eigenvectors with lowest eigenvalues may gives us insight in the way our model is well-constrained within the posterior beyond the marginalised uncertainties of single parameters.
@@
In order to check that the parameters in eigenvector space are indeed less strongly correlated, we checked the marginalised posteriors in two principal components and comfirmed that the transformed posteriors are indeed significantly closer to the idealised assumption of an elliptical distribution.
@@
The eigenvalue spectra for the posteriors of the different fits are shown in Fig. 5. We can see that they agree very well between parameter fits for different fit statistics. In order to further check the agreement of the eigenvector spaces, we considered the angles between eigenvectors for different fits. We found that the eigenvector spaces of the posteriors from $\chi^{2}$-fit and $D_{\mathrm{KL}}$-fit agree quite well, with eigenvectors with the same index corresponding to one another within angular deviations of up to 26 degrees, and less than 9 degrees if considering only indices $\lambda=0,1,2,10$. The eigenvector space of the $E[\ln P]$-fit differs more strongly from the first two eigenvector spaces. While $\mathsf{u}_{0},\,\mathsf{u}_{1},\,\mathsf{u}_{2}$ and $\mathsf{u}_{10}$ agree within 14 degrees, the eigenvectors for intermediate eigenvalues differ more strongly, with angular deviations between 17 and 54 degrees. In particular, $\mathsf{u}_{7}$ and $\mathsf{u}_{8}$ for the $E[\ln P]$-fit do not correspond well to $\mathsf{u}_{7}$ and $\mathsf{u}_{8}$ for the other two fits. This means that the posteriors agree well when it comes to the combinations of parameters subject to particularly strong or weak constraints, in particular with respect to eigenvectors $\mathsf{u}_{0},\,\mathsf{u}_{1},\,\mathsf{u}_{2}$ and $\mathsf{u}_{10}$. The intermediate eigenvectors with eigenvalues closer to unity are less certainly determined in general, so we do not expect a strong correspondence here.
@@ 5.4 Compatibility of the best model with observations
We consider now our best-fitting CAMF model, which is the $D_{\mathrm{KL}}$-fit model following our discussion in Sect. 5.2. Its parameters can be found in the fourth column of Table 4. We will discuss first the properties of the resulting model functions and then analyse how well the model matches the observed cluster sample in more depth.
@@
In Fig. 7, we plot the logarithm of the ratio of modelled and observed CAMF, both including uncertainties in age and mass as per Eqs. (7) and (14).
We observe the best agreement between CAMFs at intermediate ages and masses, where the cluster surface densities are high.
@@
In order to test the consistency of the best-fitting model with the data in more detail, we consider the CAMF in initial-mass space. We recall Eq. (10), which implies that in initial-mass space, the CAMF is given by the product of CFR and CIMF. By using the bound-mass function to infer initial cluster masses, we can thus reconstruct the CIMF and the CFR from the cluster sample data.
Figure 8: CAMF in initial mass space constructed using the bound-mass relation of the best-fitting model. Cluster densities are computed for linear bin widths in age and logarithmic bin widths in initial mass. The coloured bar on the left represents the product of model CIMF and CFR, which is constant with age. The dashed line represents the model lifetime-mass relation.
We construct the CAMF in initial-mass space in Fig. 8, using a linear age scale for bin normalisation in order to visualise the factorisation of the CAMF according to Eq. (10). As our CFR is constant and the CIMF does not depend on cluster age, this initial-mass CAMF is $\approx\beta f(M)\Delta M/\Delta\log(M/\mathrm{M}_{\odot})$ within a bin centred on initial mass $M$, independent of cluster age. We see that Fig. 8 seems compatible with the adopted constant CFR and unchanging CIMF. While we observe few higher-mass clusters with $\mu>3.5$ at young ages $\tau<8$, this can easily be explained by stochastic noise due to the generally lower number of clusters at younger ages. If we want to gain further information from this CAMF, more assumptions must be made. In particular, by assuming our model CIMF to be the correct one, we can extract the CFR from the CAMF, and vice versa to extract the CIMF.
To reconstruct the CFR, we consider the clusters with ages in an infinitesimal interval $\mathrm{d}t$ around age $t$. The number of clusters formed at that age is then $\Psi(t)\mathrm{d}t$, with a surviving fraction of
@@
The total cluster surface density at age $t$ is thus $f_{\mathrm{surv}}(t)\Psi(t)\mathrm{d}t$, and we can then reconstruct the CFR given CIMF and cluster lifetime-mass relation by computing
@@
We reconstruct the CFR from observations using the CIMF and lifetime-mass relation of our model and plot it in Fig. 9 together with the model CFR as well as a linear fit to the reconstructed CFR to check for age-dependent trends. We note that while the model CFR in principle extends to $\tau=10$, the reconstructed CFR can only be constructed up to $\tau=9.75$, beyond which no clusters were observed.
The slope of the linear fit at $-0.040\pm 0.005\mathrm{kpc^{-2}Myr^{-1}dex^{-1}}$, which is statistically significant, indicates that while the reconstructed CFR systematically deviates from a constant, it does so only slightly. As such, we interpret this as primarily a weakness of the model fit and not a fundamental incompatibility of the data with a constant CFR.
@@
For reconstructing the CIMF, we consider the total cluster surface density per unit mass $n(M)$ at a given initial mass $M$. This is given by
@@
for our model of constant CFR, where $t_{\mathrm{min}}$ is the lower age limit of the cluster sample. As such, we construct the CIMF from a cluster population given the bound-mass function to compute cluster initial masses and a constant CFR as
@@
As for the CFR, we reconstruct the CIMF using our CAMF model, in particular via bound-mass function and CFR, plotting the result in Fig. 10.
We see that the reconstructed CIMF is highly consistent with our model for intermediate masses of $\mu=2.5$ to 4. For initial masses above $\mu=4$, the reconstructed CIMF does not fall off as fast as the model CIMF, and may thus be compatible with a higher Schechter cutoff mass than present in the model. However, uncertainties here are relatively large and both CIMFs still mostly agree within the errors. At lower masses, a break in the reconstructed CIMF power law is observed. Compared to the model CIMF, a sharper transition between power laws at a higher transition mass $M_{\ast}$ also seems permissible, possibly in combination with a slightly steeper low-mass power-law slope. Modelled clusters with initial masses between $m_{\mathrm{lower}}=2\mathrm{M}_{\odot}$ and $10\mathrm{M}_{\odot}$ are not observed due to their short lifetimes, which makes it difficult to assess the CIMF below the breaking point of the high-mass power-law slope.
This also indicates that the sample (and the best model in particular) would be consistent with a higher value of the lower limit to the initial cluster masses. However, this makes little difference to the resulting CAMF as the total cluster surface density remains fixed. When setting $m_{\mathrm{lower}}=10\mathrm{M}_{\odot}$ for our best model, this results in changed parameter values $\log k/\mathrm{M}_{\odot}=-2.489$ and $\beta=0.5989\mathrm{kpc^{-2}Myr^{-1}}$. All other parameter values remain unchanged, and the mean initial cluster mass increases to $\langle M\rangle=144\mathrm{M}_{\odot}$, such that the CFR is only slightly decreased to $86.4\mathrm{M}_{\odot}\mathrm{kpc^{-2}Myr^{-1}}$.

### v2

@@ 5.2 Comparison of fit statistics
Secondly, we find that the parameters of violent relaxation are not significantly constrained beyond their priors. To quantify this, we considered the Kolmogorov-Smirnov statistic (KS-statistic) of the marginalised posteriors for the respective priors, that is, the maximum of the absolute difference between the cumulative distribution functions of the marginalised posterior and the prior. As we see in Table 6, the KS-statistic is significantly lower for $t_{\mathrm{V}}$ than for the parameters of the CIMF, independent of the fit statistic used. For $n_{\mathrm{b}}$, the KS-statistic is still below all but the lowest value found for any combination of fit statistic and CIMF parameter. The reason for this may be that while the entire cluster sample contains implicit information on the shape of the CIMF, only young clusters with $\tau<8$ contain information on violent relaxation, which is a smaller sample containing only 497 clusters. Moreover, this sample was already used to derive prior constraints, limiting the information that could be extracted additionally by the fitting procedure.
@@ 5.2.1 Construction of synthetic cluster samples
In order to better understand the differences between the fit statistics, we considered synthetic cluster samples as realisations of a given model. We constructed such a realisation by drawing a sufficient number $N$ of clusters from the model, applying random Gaussian errors according to the error model from Sect. 4.1, and then evenly distributing them across a heliocentric disc of radius $R=\sqrt{{N}/\left({\pi\Sigma_{\mathrm{tot}}}\right)}$ such that the cluster surface density corresponded to the observed value of $\Sigma_{\mathrm{tot}}=134.1\,\mathrm{kpc^{-2}}$. In analogy to the magnitude-dependent completeness limit, a completeness limit depending on the logarithm of its mass was drawn for each cluster. These limits were drawn from a Gaussian with mean and standard deviation of the form
@@
with values $w_{\mathrm{mean}}=0.60\,\mathrm{kpc},\,v_{\mathrm{mean}}=0.84\,\mathrm{kpc}$ for the mean and $w_{\mathrm{std}}=0.43\,\mathrm{kpc},\,v_{\mathrm{std}}=0.12\,\mathrm{kpc}$ for the standard deviation obtained by performing linear regression on the completeness limits of the observed cluster sample. A minimum completeness distance of 0.2 kpc was chosen to disallow divergent contributions to the synthetic CAMF. These completeness limits then also allowed the definition of a sufficient number of clusters such that the radius of the resulting heliocentric disc is at least as large at the greatest completeness distance. In practice, this was done by choosing $N$ such that $R\geq 6.8\,\mathrm{kpc}$.
The final synthetic sample was then obtained by selecting only clusters within their completeness limits. While this procedure produces samples with similar properties as the observed one by coarsely approximating the measurement and selection process used to arrive at a completeness-corrected sample, it cannot directly fix global properties of the samples such as number of clusters and total cluster surface density, which as a result vary between realisations and models.
For each of the base, $\chi^{2}$-fit, $D_{\mathrm{KL}}$-fit, and $E[\ln P]$-fit models, 5000 realisations of synthetic cluster samples were drawn and the fit statistics with respect to the corresponding models computed. While the total cluster surface density and cluster number of these realisations was not fixed, their construction causes them to reproduce the observed cluster surface density well, with a stochastic scatter of about 6 kpc-2. The corresponding number of clusters also varies, with systematic deviations from the cluster number of the observed sample being introduced by the differing structure of the model CAMFs. In particular, realisations of $\chi^{2}$-fit model and $E[\ln P]$-fit model tend to produce fewer clusters, while realisations of the $D_{\mathrm{KL}}$-fit model produce cluster numbers consistent with the observed sample.
@@ 5.2.2 Fit statistic distributions for model realisations
Further, the fit-statistic distributions allowed us assign a $Q$-value to the observed value of the fit statistic. This represents the quantile at which the observation sits and gives the fraction of model realisations with a lower degree of agreement with the model as measured by the statistic, that is, a larger value for $\chi^{2}$ and KLD, and a smaller value for $E[\ln P]$. This $Q$-value gives an indication how typical the level of disagreement between model and data would be for a realisation of the model, both in terms of atypically high or low values of the statistic. A very low $Q$-value means that it is unlikely that a random realisation of the model will disagree with the model at least as much as the data. As such, it signals that the model does not fit the data well. Conversely, a very high $Q$-value means that it is unlikely for random realisations of the model to agree with the model better than the data. In this sense, it signals that the model may be ‘overfitted’. What we refer to here as overfitting is when a model is fitted to not just a dataset’s general features, but also its specific noise. Such a model will produce spurious features, which is why this is an undesirable property. When overfitting occurs for a given combination of data and fit statistic, the data will match the model better than a typical realisation of the model itself, as a greater discrepancy is expected simply from the stochastic noise included in the forward modelling.
@@
In the $D_{\mathrm{KL}}$-fit model, the value of $\chi^{2}_{\mathrm{red}}$ falls within the central 68% (corresponding to a Gaussian $1\sigma$-interval), and the value of $D_{\mathrm{KL}}$ is still fairly typical with a $Q$-value at 14%. However, for $E[\ln P]$, we have a very low $Q$-value of 2.7%. As we discuss in Sect. 5.2.3, this may indicate the model realisations being systematically more centrally concentrated than the observed sample. The $D_{\mathrm{KL}}$-fit model samples also reproduce the observed cluster number well.
@@ 5.2.3 General remarks and conclusions
From our analysis of the fit statistics of synthetic cluster samples, it is clear that one factor which limits parameter fitting is the large impact of sampling noise on the fit statistics, which limits their ability to discriminate between models based on the observed cluster sample. This is likely because our fit statistics, as we constructed them in Sect. 4.2, 4.3 and 4.4, do not take the Poisson noise of the cluster sample into account correctly. In our sample, faint clusters have large weight, which increases their impact on the noise.
Of our best-fitting models, the $D_{\mathrm{KL}}$-fit model performs the best across all fit statistics and reproduces the cluster number when sampled. It reproduces the observed sharp age limit while still allowing for high-age high-mass clusters with $\tau>8.5,\,\mu>3.5$ when accounting for uncertainties in mass determination. A caveat remains in that the synthetic sampling of the model shows that the value of $E[\ln P]$ for data and model is lower than for typical realisations. This may suggest that the model CAMF is more concentrated in the age-mass plane than the observed CAMF. As such, we took the $D_{\mathrm{KL}}$-fit model to be the best model we found in our investigation of model parameter space.
@@ 5.3 Parameter correlations
Previously, we evaluated the uncertainties of the model parameters individually based on their marginalised one-dimensional posteriors. We next considered the full eleven-dimensional posterior to derive a better understanding of the uncertainties of the model parameters and their correlations.
As the high dimensionality of the parameter posterior distribution makes it challenging to analyse or visualise the full complexity of the posterior, we used principal component analysis in an attempt to make the posterior tractable without losing too much information. This was done by first normalising the parameters to zero mean and unit standard deviation, and then performing an eigenvector decomposition on the covariance matrix of the posterior in normalised parameter space. The resulting eigenvectors $\mathsf{u}_{i}$ for indices $i=0\ldots 10$, ordered by the size of the corresponding eigenvalues, form an orthonormal basis of the normalised parameter space and are uncorrelated in the sense that $\mathrm{Cov}(\mathsf{u}_{i},\mathsf{u}_{j})=\lambda_{i}\delta_{ij}$ where $\lambda_{i}$ is the eigenvalue corresponding to $\mathsf{u}_{i}$. This means that the eigenvalues represent the variance of the posterior distribution along the direction of the corresponding eigenvectors. If we consider the distribution in the eigenvector basis, then the $i$-th parameter has zero mean, standard deviation $\sqrt{\lambda_{i}}$ and no linear correlation to another parameter. While higher-order correlations certainly exist, this at least allows us to consider independence of these principal components as a first-order approximation. In this picture, the posterior corresponds to an ellipsoidal distribution with principal axes given by the eigenvectors $\mathsf{u}_{i}$.
As such, if $\lambda_{i}$ is smaller, then the posterior is more constrained along the axis $\mathsf{u}_{i}$, while it is less constrained if $\lambda_{i}$ is larger, with the vectors with largest eigenvalues having the greatest contribution to the uncertainty of the parameters. Conversely, the eigenvectors with lowest eigenvalues may give us insight in the way our model is well-constrained within the posterior beyond the marginalised uncertainties of single parameters.
@@
In order to check that the parameters in eigenvector space are indeed less strongly correlated, we checked the marginalised posteriors in two principal components. We confirmed that the transformed posteriors are indeed significantly closer to the idealised assumption of an elliptical distribution.
@@
The eigenvalue spectra for the posteriors of the different fits are shown in Fig. 5. We can see that they agree very well between parameter fits for different fit statistics. In order to further check the agreement of the eigenvector spaces, we considered the angles between eigenvectors for different fits. We found that the eigenvector spaces of the posteriors from $\chi^{2}$-fit and $D_{\mathrm{KL}}$-fit agree quite well, with eigenvectors with the same index corresponding to one another within angular deviations of up to 26 degrees, and less than 9 degrees if considering only indices $\lambda=0,1,2,10$. The eigenvector space of the $E[\ln P]$-fit differs more strongly from the first two eigenvector spaces. While $\mathsf{u}_{0},\,\mathsf{u}_{1},\,\mathsf{u}_{2}$ and $\mathsf{u}_{10}$ agree within 14 degrees, the eigenvectors for intermediate eigenvalues show angular deviations between 17 and 54 degrees. In particular, $\mathsf{u}_{7}$ and $\mathsf{u}_{8}$ for the $E[\ln P]$-fit do not correspond well to $\mathsf{u}_{7}$ and $\mathsf{u}_{8}$ for the other two fits. This means that the posteriors agree well when it comes to the combinations of parameters subject to particularly strong or weak constraints, in particular with respect to eigenvectors $\mathsf{u}_{0},\,\mathsf{u}_{1},\,\mathsf{u}_{2}$ and $\mathsf{u}_{10}$. The intermediate eigenvectors with eigenvalues closer to unity are less certainly determined in general, so we do not expect a strong correspondence here.
@@ 5.4 Compatibility of the best model with observations
We consider now our best-fitting CAMF model, which is the $D_{\mathrm{KL}}$-fit model following our discussion in Sect. 5.2. Its parameters can be found in the fourth column of Table 4. We discuss first the properties of the resulting model functions and then analyse how well the model matches the observed cluster sample in more depth.
@@
In Fig. 7, we plot the logarithm of the ratio of modelled and observed CAMF, both including uncertainties in age and mass as per Eqs. (7) and (14). We observe the best agreement between CAMFs at intermediate ages and masses, where the cluster surface densities are high.
@@
In order to test the consistency of the best-fitting model with the data in more detail, we considered the CAMF in initial-mass space. We recall Eq. (10), which implies that in initial-mass space, the CAMF is given by the product of CFR and CIMF. By using the bound-mass function to infer initial cluster masses, we can thus reconstruct the CIMF and the CFR from the cluster sample data.
Figure 8: CAMF in initial mass space constructed using the bound-mass relation of the best-fitting model. Cluster densities were computed for linear bin widths in age and logarithmic bin widths in initial mass. The coloured bar on the left represents the product of model CIMF and CFR, which is constant with age. The dashed line represents the model lifetime-mass relation.
We constructed the CAMF in initial-mass space in Fig. 8, using a linear age scale for bin normalisation in order to visualise the factorisation of the CAMF according to Eq. (10). As our CFR is constant and the CIMF does not depend on cluster age, this initial-mass CAMF is $\approx\beta f(M)\Delta M/\Delta\log(M/\mathrm{M}_{\odot})$ within a bin centred on initial mass $M$, independent of cluster age. We see that Fig. 8 seems compatible with the adopted constant CFR and unchanging CIMF. While we observe few higher-mass clusters with $\mu>3.5$ at young ages $\tau<8$, this can easily be explained by stochastic noise due to the generally lower number of clusters at younger ages. If we want to gain further information from this CAMF, more assumptions must be made. In particular, by assuming our model CIMF to be the correct one, we can extract the CFR from the CAMF, and vice versa to extract the CIMF.
To reconstruct the CFR, we considered the clusters with ages in an infinitesimal interval $\mathrm{d}t$ around age $t$. The number of clusters formed at that age was then $\Psi(t)\mathrm{d}t$, with a surviving fraction of
@@
The total cluster surface density at age $t$ is thus $f_{\mathrm{surv}}(t)\Psi(t)\mathrm{d}t$, and we could then reconstruct the CFR given CIMF and cluster lifetime-mass relation by computing
@@
We reconstructed the CFR from observations using the CIMF and lifetime-mass relation of our model and plot it in Fig. 9 together with the model CFR as well as a linear fit to the reconstructed CFR to check for age-dependent trends. We note that while the model CFR in principle extends to $\tau=10$, the reconstructed CFR could only be constructed up to $\tau=9.75$, beyond which no clusters were observed. The slope of the linear fit at $-0.040\pm 0.005\mathrm{kpc^{-2}Myr^{-1}dex^{-1}}$, which is statistically significant, indicates that while the reconstructed CFR systematically deviates from a constant, it does so only slightly. As such, we interpret this as primarily a weakness of the model fit and not a fundamental incompatibility of the data with a constant CFR.
@@
For reconstructing the CIMF, we considered the total cluster surface density per unit mass $n(M)$ at a given initial mass $M$. This is given by
@@
for our model of constant CFR, where $t_{\mathrm{min}}$ is the lower age limit of the cluster sample. As such, we constructed the CIMF from a cluster population given the bound-mass function to compute cluster initial masses and a constant CFR as
@@
As for the CFR, we reconstructed the CIMF using our CAMF model, in particular via bound-mass function and CFR, plotting the result in Fig. 10. We see that the reconstructed CIMF is highly consistent with our model for intermediate masses of $\mu=2.5$ to 4. For initial masses above $\mu=4$, the reconstructed CIMF does not fall off as fast as the model CIMF, and may thus be compatible with a higher Schechter cutoff mass than present in the model. However, uncertainties here are relatively large and both CIMFs still mostly agree within the errors. At lower masses, a break in the reconstructed CIMF power law is observed. Compared to the model CIMF, a sharper transition between power laws at a higher transition mass $M_{\ast}$ also seems permissible, possibly in combination with a slightly steeper low-mass power-law slope. Modelled clusters with initial masses between $m_{\mathrm{lower}}=2\mathrm{M}_{\odot}$ and $10\mathrm{M}_{\odot}$ are not observed due to their short lifetimes, which makes it difficult to assess the CIMF below the breaking point of the high-mass power-law slope.
This also indicates that the sample (and the best model in particular) would be consistent with a higher value of the lower limit to the initial cluster masses. However, this makes little difference to the resulting CAMF as the total cluster surface density remains fixed. When setting $m_{\mathrm{lower}}=10\mathrm{M}_{\odot}$ for our best model, this resulted in changed parameter values $\log k/\mathrm{M}_{\odot}=-2.489$ and $\beta=0.5989\mathrm{kpc^{-2}Myr^{-1}}$. All other parameter values remained unchanged, and the mean initial cluster mass increased to $\langle M\rangle=144\mathrm{M}_{\odot}$, such that the CFR was only slightly decreased to $86.4\mathrm{M}_{\odot}\mathrm{kpc^{-2}Myr^{-1}}$.